OpenStudy (anonymous):

At the last football game, 4 out of every 9 students wore school colors. Based on this information, if 450 students attend the next football game, how many students will wear school colors? A) 150 B) 175 C) 200 D) 225

OpenStudy (anonymous):

4/9 = x/450 Cross multiply 4(450) = 9x 1800 = 9x /9 /9 200 = x 200 is the answer. C

OpenStudy (anonymous):

thx can u help with a few more?

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

One gram of fat contains 180% of the calories of a gram of a carbohydrate. How many times more calories are there in 8 grams of fat than there are in 8 grams of carbohydrate? A) 0.23 B) 1.44 C) 14.4 D) 23.3

OpenStudy (anonymous):

I don't think the answer is actually there.

OpenStudy (anonymous):

0-0

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

C. 1.8 (180%) * 8 grams=14.4

OpenStudy (anonymous):

Ben believes that most of the automobiles in his neighborhood are SUVs. He does a house-to-house survey of parked cars. He surveys 45 cars and only 8 of the cars he finds are SUVs. Yet fully 50% of the town’s automobiles are SUVs. What is the MOST LIKELY explanation for these two facts? A) Ben’s sample is too small to be representative of the entire town. B) SUVs are sometimes hard to distinguish from non-SUVs and Ben made errors. C) Most of the SUVs were on the road, and less popular models were left at home. D) Ben only looked at parked cars instead of asking residents what kind of car they drive.

OpenStudy (anonymous):

That's not the same as how many times more calories.

OpenStudy (anonymous):

This is kind of a BS question, but D maybe.